1. During the experiment, Althea prepares a specific chemical to test on. The chemical has 56% C, 18%H, 32% O. a. Predict the Empirical formula b. If the amount of that compound being used is 120 g/mole, what will be its molecular formula? *
2. Marco uses Ammonia (NH3) on his titration experiment. If the experimental usage of the NH3 is 68g/mole. What is the Molecular formula of of NH3 being used? *
3. Determine the empirical formulas for compounds with the following percent compositions: (a) 15.8% carbon and 84.2% sulfur (b) 40.0% carbon, 6.7% hydrogen, and 53.3% oxygen *
1
Expert's answer
2020-10-16T08:24:38-0400
  1. (a) Moles of C = "=\dfrac{56}{12}=4.67" ( mole ratio= 4.67/2 =2)


Moles of H"=\dfrac{18}{1} = 18" ( mole ratio = 18/2 =9)

Moles of O "=\dfrac{32}{16}=2" ( mole ratio = 2/2 = 1)


EF is therefore; C2H9O1

= C2H9O


(b) Molecular formula = Empirical formula x n


"n=\dfrac{Molecular formular mass}{empirical formula mass} = \dfrac{120}{(24+9+16)}" = 2.4 (approx. 2)


MF = (C2H9O1) x 2


= C4H18O2


2.

MF = EF x n

"n= \dfrac{68}{(14+3)} = 4"


MF =(NH3) x 4


= N4H13


3.

(a) Moles of C "=\dfrac{15.8}{12} = 1.317" ( Mole ratio = 1.317/1.317 = 1)


Moles of S "= \dfrac{84.2}{32.1} = 2.623" (mole ratio = 2.623/1.317 = 2)


EF = C1S2


= CS2


(b) Moles of C "=\dfrac{40}{12} = 3.333" (mole ratio = 3.333/3.331 = 1)


Moles of H "=\dfrac{6.7}{1} = 6.7" (mole ratio = 6.7/3.331 = 2)


Moles of O "=\dfrac{53.3}{16} = 3.331" (Mole ratio = 3.331/3.331 = 1)


EF = C1H2O1


= CH2O